OpenStudy (anonymous):

Use Part 1 of the Fundamental Theorem of Calculus to find the derivative of the function. g(s)= integral 2 to s (t-t^3)^5

OpenStudy (anonymous):

the derivative of the integral is the integrand replace \(t\) by \(s\) in the integrand and you are done

OpenStudy (anonymous):

this is much much easier than it seems

OpenStudy (anonymous):

The fundamental theorem basically says that when you have an integral going from a number to some variable, we substitute that variable for the inside and multiply by it's derivative.

OpenStudy (anonymous):

In this case the derivative of s is just 1 so we can disregard that.

OpenStudy (anonymous):

for example, the derivative of \[\int_2^x\frac{t^2+2}{t-1}dt\] is \[=\frac{x^2+2}{x-1}\]

OpenStudy (anonymous):

^ Yeah that's good.

OpenStudy (anonymous):

thanks...d

OpenStudy (anonymous):

No problem :) .